# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 2. Preliminaries
We collect some notation and basic concepts which will be used in this paper. We denote by ${\boldsymbol{x}}$ (resp., ${\boldsymbol{y}}$) the $n$-tuple (resp., $m$-tuple) of variables $(x_{1},\ldots,x_{n})$ (resp., $(y_{1},\ldots,y_{m})$). The symbol $\mathbb{N}$ (resp., ${\mathbb{R}}$, ${\mathbb{R}}_{+}$) denotes the set of nonnegative integers (resp., real numbers, nonnegative real numbers). For positive integer $n\in\mathbb{N}$, denote by $[n]$ the set $\{1,\ldots,n\}$. Denote by ${\mathbb{R}}^{p}$ (resp. ${\mathbb{R}}^{l_{1}\times l_{2}}$, ${\mathbb{S}}^{p}$, ${\mathbb{S}}_{+}^{p}$) the $p$-dimensional real vector (resp. $l_{1}\times l_{2}$ real matrix, $p\times p$ symmetric real matrix, $p\times p$ PSD matrix) space. Denote by +${\mathbb{R}}^{n}_{+}$ the nonnegative orthant of ${\mathbb{R}}^{n}$. For ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ (resp., $N\in{\mathbb{R}}^{l_{1}\times l_{2}}$), the symbol ${\boldsymbol{v}}^{\intercal}$ (resp., $N^{\intercal}$) denotes the transpose of ${\boldsymbol{v}}$ (resp., $N$). For a matrix $N\in{\mathbb{R}}^{p\times p}$, $\mbox{tr}\left(N\right)$ denotes its trace. For two matrices $N_{1}$ and $N_{2}$, $N_{1}\otimes N_{2}$ denotes the Kronecker product of $N_{1}$ and $N_{2}$. For two matrices $N_{1}$ and $N_{2}$ of the same size, $\langle N_{1},N_{2}\rangle$ denotes the inner product $\mbox{tr}\left(N_{1}^{\intercal}N_{2}\right)$ of $N_{1}$ and $N_{2}$. The notation $I_{p}$ denotes the $p\times p$ identity matrix. For any $t\in{\mathbb{R}}$, $\lceil t\rceil$ (resp., $\lfloor t\rfloor$) denotes the smallest (resp., largest) integer that is not smaller (resp., larger) than $t$. For ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$, $\|{\boldsymbol{u}}\|$ denotes the standard Euclidean norm of ${\boldsymbol{u}}$. For $N\in{\mathbb{R}}^{l_{1}\times l_{2}}$, $\|N\|$ denotes the spectral norm of $N$. For a vector ${\boldsymbol{\alpha}}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}$, let $|{\boldsymbol{\alpha}}|=\alpha_{1}+\cdots+\alpha_{n}$. For a set $A$, we use $|A|$ to denote its cardinality. For $k\in\mathbb{N}$, let $\mathbb{N}^{n}_{k}\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid|{\boldsymbol{\alpha}}|\leq k\}$ and $|\mathbb{N}^{n}_{k}|=\binom{n+k}{k}$ be its cardinality. For variables ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$, ${\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ denotes the monomial $x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. Let ${\mathbb{R}}[{\boldsymbol{x}}]$ (resp. ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$) denote the set of real polynomials (resp. $p\times p$ symmetric real polynomial matrices) in ${\boldsymbol{x}}$. For $h\in{\mathbb{R}}[{\boldsymbol{x}}]$ (resp., $h\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]$), we denote by $\deg(h)$ (resp., $\deg_{{\boldsymbol{x}}}(h)$) its total degree in ${\boldsymbol{x}}$. For a polynomial matrix $T({\boldsymbol{x}})=[T_{ij}({\boldsymbol{x}})]$ (resp., $T({\boldsymbol{x}},{\boldsymbol{y}})=[T_{ij}({\boldsymbol{x}},{\boldsymbol{y}})]$), denote $\deg(T)\coloneqq\max_{i,j}\deg(T_{ij})$ (resp., $\deg_{{\boldsymbol{x}}}(T)\coloneqq\max_{i,j}\deg_{{\boldsymbol{x}}}(T_{ij})$). For $k\in\mathbb{N}$, denote by ${\mathbb{R}}[{\boldsymbol{x}}]_{k}$ (resp., ${\mathbb{S}}[{\boldsymbol{x}}]^{p}_{k}$) the subset of ${\mathbb{R}}[{\boldsymbol{x}}]$ (resp., ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$) of degree up to $k$. For any $P({\boldsymbol{x}})=[P_{ij}({\boldsymbol{x}})]\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $Q({\boldsymbol{x}},{\boldsymbol{y}})=[Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})]\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$, denote
@@
2.1. SOS matrices and positivstellensätz for polynomial matrices
For a polynomial $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$, if there exist polynomials $f_{1}({\boldsymbol{x}}),\ldots,f_{t}({\boldsymbol{x}})$ such that $f({\boldsymbol{x}})=\sum_{i=1}^{t}f_{i}({\boldsymbol{x}})^{2}$, then we call $f({\boldsymbol{x}})$ an SOS. A polynomial matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is said to be an SOS matrix if there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. For $d\in\mathbb{N}$, denote by $[{\boldsymbol{x}}]_{d}$ the canonical basis of ${\mathbb{R}}[{\boldsymbol{x}}]_{d}$, i.e.,
$[{\boldsymbol{x}}]_{d}\coloneqq[1,\ x_{1},\ x_{2},\ \cdots,\ x_{n},\ x_{1}^{2},\ x_{1}x_{2},\ \cdots,\ x_{n}^{d}]^{\intercal},$ (2)
whose cardinality is $|\mathbb{N}^{n}_{d}|=\binom{n+d}{d}$. With $d=\deg(T)$, we can write $T({\boldsymbol{x}})$ as
@@ 2.1. SOS matrices and positivstellensätz for polynomial matrices
where $Q$ is the vector of coefficient matrices of $T({\boldsymbol{x}})$ with respect to $[{\boldsymbol{x}}]_{d}$. Hence, $\Sigma({\boldsymbol{x}})$ is an SOS matrix with respect to $[{\boldsymbol{x}}]_{d}$ if there exists some $Q\in{\mathbb{R}}^{l\times p|\mathbb{N}^{n}_{d}|}$ satisfying
$\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}(Q^{\intercal}Q)([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
We thus have the following results.
@@
[45, Lemma 1] A polynomial matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is an SOS matrix with respect to the monomial basis $[{\boldsymbol{x}}]_{d}$ if and only if there exists $Z\in\mathbb{S}_{+}^{p|\mathbb{N}^{n}_{d}|}$ such that
$\Sigma({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}Z([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
@@
Let $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be an SOS matrix and $\Sigma_{k}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{k\times k}$ be a principal submatrix of $\Sigma({\boldsymbol{x}})$ whose rows and columns are indexed by $(p_{1},\ldots,p_{k})$ with $1\leq p_{1}<\cdots<p_{k}\leq p$, then $\Sigma_{k}({\boldsymbol{x}})$ is an SOS matrix.
@@
As $\Sigma({\boldsymbol{x}})$ is an SOS matrix, there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. Denote by $T_{k}({\boldsymbol{x}})$ the submatrix of $T({\boldsymbol{x}})$ consisting of the colums of $T(x)$indexed by $(p_{1},\ldots,p_{k})$. Then, $\Sigma_{k}({\boldsymbol{x}})=T_{k}({\boldsymbol{x}})^{\intercal}T_{k}({\boldsymbol{x}})$ and hence is an SOS matrix. ∎
@@
When $p=1$, the product $\left\langle A,B\right\rangle_{1}$ coicides with the matrix inner product $\langle A,B\rangle=\mbox{tr}\left(A^{\intercal}B\right)$.
@@
Let $H({\boldsymbol{x}})=\mbox{diag}(H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))$ be a block diagonal matrix, then for any SOS matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr}$ where $p\in\mathbb{N}$ and $r=r_{1}+\cdots+r_{t}$, there are SOS matrices $\Sigma_{j}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$, $j\in[t]$, such that
$\left\langle\Sigma({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
@@
For each $j\in[t]$, let $\Sigma_{j}({\boldsymbol{x}})$ be the $pr_{j}\times pr_{j}$ principal submatrix of $\Sigma({\boldsymbol{x}})$ whose rows and columns are indexed by
@@
As $H({\boldsymbol{x}})$ is block diagonal, by the definition of the mapping .,.$\left\langle.,.\right\rangle_{p}$, it is easy to see that
$\left\langle\Sigma({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
By Lemma 2.1, each $\Sigma_{j}({\boldsymbol{x}})$ is an SOS matrix. ∎
The matrix quadratic module $\mathcal{Q}^{p}(\mathbf{H})$ generated by $\mathbf{H}$ is defined as
$\mathcal{Q}^{p}(\mathbf{H})\coloneqq\left\{\Sigma_{0}({\boldsymbol{x}})+\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr_{j}},j\in[t],\ \text{are SOS}\right\}.$
@@
(Scherer-Hol’s Positivestellensatz) Let Assumption 2.1 and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be PD on $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$. Then, $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{H})$.
@@
It was proved in [45, Corollary 1] for the case $t=1$. The case $t>1$ can be derived from [45, Corollary 1] and Lemma 2.2. ∎
2.2. Matrix-valued measures and moment problem
Now we recall some background on the concept of matrix-valued measures. Denote by $\mathfrak{B}(\mathcal{X})$ the smallest $\sigma$-algebra generated from the open subsets of $\mathcal{X}$ and by $\mathfrak{m}(\mathcal{X})$ the set of all finite Borel measures on $\mathcal{X}$. The support $\mathrm{supp}(\phi)$ of a Borel measure $\phi\in\mathfrak{m}(\mathcal{X})$ is the (unique) smallest closed set $\mathbf{A}\in\mathfrak{B}(\mathcal{X})$ such that $\phi(\mathcal{X}\setminus\mathbf{A})=0$. Let $\phi_{ij}\in\mathfrak{m}(\mathcal{X})$, $i,j=1,\ldots,p$. The $p\times p$ matrix-valued measure $\Phi$ on $\mathcal{X}$ is defined as the matrix-valued function $\Phi\colon\mathfrak{B}(\mathcal{X})\to{\mathbb{R}}^{p\times p}$ with
$\Phi(\mathbf{A})\coloneqq[\phi_{ij}(\mathbf{A})]\in{\mathbb{R}}^{p\times p},\quad\forall\mathbf{A}\in\mathfrak{B}(\mathcal{X}).$
If $\phi_{ij}=\phi_{ji}$ for all $i,j=1,\ldots,p$, we call $\Phi$ a symmetric matrix-valued measure. If ${\boldsymbol{v}}^{\intercal}\Phi(\mathbf{A}){\boldsymbol{v}}\geq 0$ holds for all $\mathbf{A}\in\mathfrak{B}(\mathcal{X})$ and for all column vectors ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$, we call $\Phi$ a PSD matrix-valued measure. The set $\mathrm{supp}(\Phi)\coloneqq\bigcup_{i,j=1}^{p}\mathrm{supp}(\phi_{ij})$ is called the support of the matrix-valued measure $\Phi$. We denote by $\mathfrak{M}^{p}_{+}(\mathcal{X})$ the set of all $p\times p$ PSD symmetric matrix-valued measures on $\mathcal{X}$.
For a polynomial matrix $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and a matrix-valued measure $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$, the integral of $H$ with respect to $\Phi$ is defined by
$\int_{\mathcal{X}}H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\coloneqq\int_{\mathcal{X}}\ \mbox{tr}\left(H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\right)=\sum_{i,j}\int_{\mathcal{X}}\ H_{ij}({\boldsymbol{x}})\mathrm{d}\phi_{ij}({\boldsymbol{x}}).$
Let $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ be a multi-indexed sequence of symmetric matrices in ${\mathbb{S}}^{p}$. We define a linear functional $\mathscr{L}_{\mathbf{S}}\colon{\mathbb{S}}[{\boldsymbol{x}}]^{p}\to{\mathbb{R}}$ in the following way:
$\mathscr{L}_{\mathbf{S}}(H)\coloneqq\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}\mbox{tr}\left(H_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\alpha}}}\right),\quad\forall H({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}H_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}.$
We call $\mathscr{L}_{\mathbf{S}}$ the Riesz functional associated to the sequence $\mathbf{S}$. We say the sequence $\mathbf{S}$ has a matrix-valued representing measure $\Phi=[\phi_{ij}]\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ if
$S_{{\boldsymbol{\alpha}}}=\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})\coloneqq\left[\int_{\mathbf{K}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi_{ij}({\boldsymbol{x}})\right]_{i,j\in[p]},\ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}.$ (3)
The following theorem is a matrix version of Haviland’s theorem ([8, 46]).
Theorem 2.2 (Haviland’s theorem for polynomial matrices).
[8, Theorem 3] A given sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ has a matrix-valued representing measure $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ if and only if $\mathscr{L}_{\mathbf{S}}(H)\geq 0$ for all $H({\boldsymbol{x}})$ which are PSD on $\mathcal{X}$.
[…]
By a $\star$-algebra we mean an algebra $\mathcal{A}$ over the field ${\mathbb{K}}={\mathbb{R}}$ or ${\mathbb{K}}={\mathbb{C}}$ equipped with an involution $a\to a^{\star}$ of $\mathcal{A}$ into itself such that $(\lambda a+\mu b)^{\star}=\bar{\lambda}a^{\star}+\bar{\mu}b^{\star}$, $(ab)^{\star}=b^{\star}a^{\star}$ and $(a^{\star})^{\star}$ for all $a,b\in\mathcal{A}$ and $\lambda,\mu\in{\mathbb{K}}$. Denote by $1$ the unit element of $\mathcal{A}$. If $a$ is an operator on a Hilbert space, denote by $\mathcal{D}(a)$ its domain.
[…]
A state of $\mathcal{A}$ is a $\star$-linear functional $\Pi$ on $\mathcal{A}$ such that $\Pi(1)=1$ and $\Pi(a^{\star}a)\geq 0$ for all $a\in\mathcal{A}$. If $\pi$ is a $\star$-representation of $\mathcal{A}$ and $\varphi\in\mathcal{D}(\pi)$ is a unit vector, then $\Pi(\cdot)\coloneqq\langle\pi(\cdot)\varphi,\varphi\rangle$ is a state of $\mathcal{A}$. Conversely, for each state $\Pi$ on $\mathcal{A}$ there exist a $\star$-representation of $\mathcal{A}$, called the GNS representation $\pi_{\Pi}$ of $\Pi$, and a vector $\varphi_{\Pi}\in\mathcal{D}(\pi_{\Pi})$ such that $\mathcal{D}(\pi_{\Pi})=\pi_{\Pi}(\mathcal{A})\varphi_{\Pi}$ and $\Pi(a)=\langle\pi_{\Pi}(a)\varphi_{\Pi},\varphi_{\Pi}\rangle$ for all $a\in\mathcal{A}$.

### v4

@@ 2. Preliminaries
[…]
@@
Let us recall some background on SOS matrices and positivstellensätz for polynomial matrices. For a polynomial $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$, if there exist polynomials $f_{1}({\boldsymbol{x}}),\ldots,f_{t}({\boldsymbol{x}})$ such that $f({\boldsymbol{x}})=\sum_{i=1}^{t}f_{i}({\boldsymbol{x}})^{2}$, then we call $f({\boldsymbol{x}})$ an SOS. A polynomial matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is said to be an SOS matrix if there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. For $d\in\mathbb{N}$, denote by ${\boldsymbol{x}}]_{d}$ the canonical basis of ${\mathbb{R}}[{\boldsymbol{x}}]_{d}$, i.e.,
${\boldsymbol{x}}]_{d}\coloneqq[1,\ x_{1},\ x_{2},\ \cdots,\ x_{n},\ x_{1}^{2},\ x_{1}x_{2},\ \cdots,\ x_{n}^{d}]^{\intercal},$ (2)
whose cardinality is $|\mathbb{N}^{n}_{d}|=\binom{n+d}{d}$. With $d=\deg T$, we can write $T({\boldsymbol{x}})$ as
@@
where $Q$ is the vector of the coefficient matrices of $T({\boldsymbol{x}})$ with respect to ${\boldsymbol{x}}]_{d}$. Hence, $S({\boldsymbol{x}})$ is an SOS matrix with respect to ${\boldsymbol{x}}]_{d}$ if there exists some $Q\in{\mathbb{R}}^{l\times p|\mathbb{N}^{n}_{d}|}$ satisfying
$S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}(Q^{\intercal}Q)([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
Thus, we have the following results.
@@
[42, Lemma 1] A polynomial matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is an SOS matrix with respect to the monomial basis ${\boldsymbol{x}}]_{d}$ if and only if there exists $Z\in\mathbb{S}_{+}^{p|\mathbb{N}^{n}_{d}|}$ such that $S({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}Z([{\boldsymbol{x}}]_{d}\otimes I_{p})$.
@@
Let $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be an SOS matrix and $S_{k}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{k\times k}$ be a principal submatrix of $S({\boldsymbol{x}})$ whose rows and columns are indexed by $(p_{1},\ldots,p_{k})$ with $1\leq p_{1}<\cdots<p_{k}\leq p$, then $S_{k}({\boldsymbol{x}})$ is an SOS matrix.
@@
As $S({\boldsymbol{x}})$ is an SOS matrix, there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. Denote by $T_{k}({\boldsymbol{x}})$ the submatrix of $T({\boldsymbol{x}})$ consisting of the columns of T(x)T(x)indexed by $(p_{1},\ldots,p_{k})$. Then $S_{k}({\boldsymbol{x}})=T_{k}({\boldsymbol{x}})^{\intercal}T_{k}({\boldsymbol{x}})$ and hence is an SOS matrix. ∎
@@
When $p=1$, the product $\left\langle A,B\right\rangle_{1}$ coincides with the matrix inner product $\langle A,B\rangle=\mbox{tr}\left(A^{\intercal}B\right)$.
@@
Let $H({\boldsymbol{x}})=\mbox{diag}(H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))$ be a block diagonal matrix, then for any SOS matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr}$ where $p\in\mathbb{N}$ and $r=r_{1}+\cdots+r_{t}$, there are SOS matrices $S_{j}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$, $j\in[t]$, such that
$\left\langle S({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
@@
For each $j\in[t]$, let $S_{j}({\boldsymbol{x}})$ be the $pr_{j}\times pr_{j}$ principal submatrix of $S({\boldsymbol{x}})$ whose rows and columns are indexed by
@@
As $H({\boldsymbol{x}})$ is block diagonal, by the definition of the mapping $\left\langle.,.\right\rangle_{p}$, it is easy to see that
$\left\langle S({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
By Lemma 2.1, each $S_{j}({\boldsymbol{x}})$ is an SOS matrix. ∎
The matrix quadratic module $\mathcal{Q}^{p}(\mathbf{H})$ generated by $\mathbf{H}$ is defined as the cone
$\mathcal{Q}^{p}(\mathbf{H})\coloneqq\left\{S_{0}({\boldsymbol{x}})+\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr_{j}},j\in[t],\ \text{are SOS}\right\}.$
The $k$-th truncated matrix quadratic module $\mathcal{Q}^{p}_{k}(\mathbf{H})$ is a subcone of $\mathcal{Q}^{p}(\mathbf{H})$ in which the degrees of the polynomial matrices $S_{0}({\boldsymbol{x}})$ and $\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}$ are bounded by $2k$.
@@
(Scherer-Hol’s Positivestellensatz) Let Assumption 2.1 and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be PD on $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$. Then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{H})$.
@@
It was proved in [42, Corollary 1] for the case $t=1$. The case $t>1$ can be derived from [42, Corollary 1] and Lemma 2.2. ∎